# Machine Learning Nonadiabatic Dynamics: Eliminating Phase Freedom of Nonadiabatic Couplings with the State-Interaction State-Averaged Spin-Restricted Ensemble-Referenced Kohn–Sham Approach

### v1

@@ Abstract
Excited-state molecular dynamics (ESMD) simulations near conical intersections (CIs) pose significant challenges when using machine learning potentials (MLPs). Although MLPs have gained recognition for their integration into mixed quantum-classical (MQC) methods, such as trajectory surface hopping (TSH), and their capacity to model correlated electron-nuclear dynamics efficiently, difficulties persist in managing nonadiabatic dynamics. Specifically, singularities at CIs and double-valued coupling elements result in discontinuities that disrupt the smoothness of predictive functions. Partial solutions have been provided by learning diabatic Hamiltonians with phaseless loss functions to these challenges. However, a definitive method for addressing the discontinuities caused by CIs and double-valued coupling elements has yet to be developed. Here, we introduce the phaseless coupling term, $\Delta^{2}$, derived from the square of the off-diagonal elements of the diabatic Hamiltonian in the SSR(2,2) formalism. This approach improves the stability and accuracy of the MLP model by addressing the issues arising from CI singularities and double-valued coupling functions. We apply this method to the penta-2,4-dieniminium cation (PSB3), demonstrating its effectiveness in improving MLP training for ML-based nonadiabatic dynamics. Our results show that the $\Delta^{2}$ based ML-ESMD method can reproduce ab initio ESMD simulations, underscoring its potential and efficiency for broader applications, particularly in large-scale and long-timescale ESMD simulations.
@@ 1 Introduction
From a practical perspective, mixed quantum-classical (MQC) approaches17, 18, 19, 20, 21 are commonly used where MQC methods treat electronic propagation quantum mechanically while the nuclei move classically along multiple trajectories. Since conventional photochemical reactions occur over timescales ranging from hundreds of femtoseconds to nanoseconds, a large number of electronic structure calculations are required during an ESMD simulation. For instance, predicting PESs for 100,000 nuclear configurations is necessary to simulate dynamics for 1 ps with a time step size of 1 fs and 100 nuclear trajectories. Moreover, ESMD is highly sensitive to the accuracy of nonadiabatic coupling vectors (NACVs), particularly in cases involving conical intersections (CIs). In such scenarios, quantum chemical methods that adequately account for electronic correlation must be chosen. Although there are highly accurate quantum mechanical methods, e.g. equation-of-motion coupled cluster singles and doubles (EOM-CCSD)22, 23, MQC simulations require hundreds or even thousands of trajectories, making it computationally prohibitive to use such high-level ab initio theories, even for small systems. This makes the trade-off between accuracy and computational cost a central issue in MQC dynamics, presenting an ongoing challenge in balancing accuracy with feasibility.
@@
Over the past decade, machine learning potentials (MLPs) have enabled us to overcome the accuracy-cost trade-off in theoretical chemistry. MLPs can predict molecular properties with an accuracy comparable to traditional ab initio theories but at a significantly lower computational cost. 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46 These advances in ML-based property prediction can be integrated into the MQC dynamics by replacing quantum mechanical calculations with MLPs. However, applying MLPs to ESMD introduces unique challenges that are not typically encountered in ground state MD simulations. A key difficulty is the prediction of NACVs near CIs of PESs. The NACV formula incorporates the energy difference between two electronic states in the denominator. Even minor energy errors can cause significant NACV inaccuracies near CIs. Additionally, NACVs are not differentiable and diverge at CIs. Various studies have proposed solutions to the singularities at CIs. The first approach is to avoid the direct prediction of NACV by using a method that does not require NACV. The simplicity of the Landau–Zener formalism47 has attracted researchers and has been employed in various studies48, 49. The Zhu-Nakamura method50, based on the Landau–Zener formalism, has also been utilized in ML-based noadiabatic dyanmics51, 52. Another approach focused on learning diabatic Hamiltonian instead of adiabatic Hamiltonian53, 54, 55. This provided smooth and differentiable predictions for energy and NACVs near CIs.
@@
While the issue of CI singularity can be addressed through diabatic transformation, another challenge arises from the double-valued nature of the coupling elements. This double-valued nature, known as the geometric phase effect56, imparts a similar double-valued character to NACVs. As a result, NACVs derived from electronic structure calculations have arbitrary signs, leading to discontinuities that hinder efficient model training. To circumvent this problem, a phaseless loss function was introduced57 and applied in several authors of this work58. However, this approach does not fully resolve the problem, as the training set contains two values (one positive and one negative) for the same molecular geometry. A more recent work proposed an auxiliary single-valued function59, where the double-valued function, $\Gamma$, to eliminate the double-value problem. However, this method faces practical limitations, as the size of $\Gamma$ increases with system size, making it unsuitable for larger systems.
@@
Here, we introduce a phaseless SSR(2,2)60, 61, 62 formalism to address both the CI singularities and the double value problem of coupling elements. Our approach is based on deriving the SSR(2,2) equations using the phaseless coupling, $\Delta^{2}$, which represents the square of the off-diagonal elements in the diabatic Hamiltonian, rather than the phase-dependent $\Delta$. We trained equivariant message-passing MLPs to model diabatic PESs and the phaseless $\Delta^{2}$ for the penta-2,4-dieniminium cation (PSB3). This method resolves the phase ambiguity of coupling elements and enhances the accuracy of MLP prediction of NACV in SSR(2,2) ESMD simulations.
@@ 2 Method
2.1 Phaseless SSR(2,2) formalism
We employ the SSR(2,2) methodology to calculate excited electronic states. In the SSR(2,2) method, an ensemble of ground and excited microstates provides $(2\times 2)$ diabatic Hamiltonian, incorporating both perfectly spin-paired singlet (PPS) and open-shell singlet (OSS) configurations. These configurations are determined by two Kohn-Sham (KS) orbitals, $\phi_{a}$ and $\phi_{b}$, along with their fractional occupation numbers (FONs), $n_{a}$ and $n_{b}$. From these orbitals, a total of six microstates can be configured, enabling the construction of both PPS and OSS configurations as
@@ 2.1 Phaseless SSR(2,2) formalism
$f(n_{a},n_{b})=(n_{a}n_{b})^{(1-(1/2)((n_{a}n_{b}+\delta)/(1+\delta)))}$ with $\delta=$0.4. State-averaged energy functional $E^{SA}$ is constructed from $E^{PPS}$ and $E^{OSS}$, by $E^{SA}=w_{PPS}E^{PPS}+w_{OSS}E^{OSS}$. Here, we choose equiensemble condition, $w_{PPS}=w_{PPS}=1/2$. From the optimization of orbitals, we obtain the pseudo-Fock equation,
@@
$\hat{F}_{p}=\sum\limits_{L}C^{SA}_{L}{\frac{n^{L}_{p\alpha}\hat{F}^{L}_{\alpha}+n^{L}_{p\beta}\hat{F}^{L}_{\beta}}{n_{p}}}$ (4)
Here, $n_{p}$ is an averaged occupation $n_{p}=\sum\limits_{L}C_{L}^{SA}(n^{L}_{p\alpha}+n^{L}_{p\beta})$, $C_{L}^{SA}$ is an averaged weight $C_{L}^{SA}=w_{PPS}C_{L}^{PPS}+w_{OSS}C_{L}^{OSS}$, and $\epsilon^{SA}_{pq}$ is a Lagrange matrix element. Finally, we can construct $(2\times 2)$ secular equation to obtain ground and first excited state energies, $E_{-}^{SSR}$ and $E_{+}^{SSR}$, as
@@
The state interaction between PPS and OSS configuration, $\Delta^{SA}$, is calculated from a Lagrange matrix element and FONs as
$\Delta^{SA}=(\sqrt{n_{a}}-\sqrt{n_{b}})\epsilon^{SA}_{pq}$ (6)
In the SSR(2,2) method, the NACV between two adiabatic states can be directly calculated from the diabatic elements $E^{PPS}$,$E^{OSS}$, and $\Delta^{SA}$ by
$d_{01}=\frac{1}{E_{+}^{SSR}-E_{-}^{SSR}}((a_{00}a_{01}-a_{10}a_{11})\nabla(E^{PPS}-E^{OSS})+(a_{00}a_{11}-a_{01}a_{10})\nabla\Delta^{SA})$ (7)
The coefficients $a_{ij}$ can be expressed with $E^{PPS}$, $E^{OSS}$, and $\Delta^{SA}$ by solving the secular equation. Our key observation is that the NACV can be rewritten based on the phaseless quantity $\Delta^{2}$ as follows:
$d_{01}=\frac{nm}{E_{+}^{SSR}-E_{-}^{SSR}}((-2\Delta^{2})\nabla(E^{PPS}-E^{OSS})+\frac{1}{2}(E^{OSS}-E^{PPS})\nabla\Delta^{2})$ (8)
@@
$n=1/\sqrt{(E_{+}-E^{OSS})^{2}+\Delta^{2}}$ and $m=1/\sqrt{(E_{-}-E^{OSS})^{2}+\Delta^{2}}$, respectively. The significance of Eq. 8 lies in the complete substitution of $\Delta$ with $\Delta^{2}$. By this substitution, we no longer suffer from the phase arbitrariness that is critical for building MLPs for ESMD. We note that the adiabatic state energy and gradient can be rewritten by $\Delta^{2}$ instead of $\Delta$ as
@@
$\nabla E_{+}^{SSR}$ $=$ $m^{2}((E_{-}^{SSR}-E^{OSS})^{2}\nabla E^{PPS}+\Delta^{2}\nabla E^{OSS}+(E_{-}^{SSR}-E^{OSS})\nabla\Delta^{2})$
$\nabla E_{-}^{SSR}$ $=$ $n^{2}((E_{+}^{SSR}-E^{OSS})^{2}\nabla E^{PPS}+\Delta^{2}\nabla E^{OSS}+(E_{+}^{SSR}-E^{OSS})\nabla\Delta^{2})$ (10)
2.2 Machine learning potential for phaseless SSR(2,2)
@@
We employed the equivariant message-passing MLP, NequIP29, implemented in the Bayesian Atoms Modeling (BAM) package, to train three key properties: $E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$. Three independent MLP models were trained to predict each property. The analytical gradients of $\Delta^{2}$, denoted as $\nabla\Delta^{2}$, were computed by expanding $\Delta^{2}$ as $2\Delta\nabla\Delta$. To ensure that $\Delta^{2}$ remained positive, we applied inductive bias layers, such as ReLU activation function, enforcing a non-negative value constraint.
[…]
@@
$\mathcal{L}=\frac{\lambda_{E}}{B}\sum_{b=1}^{B}\left(\hat{E}_{b}-E_{b}\right)^{2}+\frac{\lambda_{F}}{3BN}\sum_{i=1}^{B\cdot N}\sum_{\alpha=1}^{3}\left(-\frac{\partial\hat{E}}{\partial r_{i,\alpha}}-F_{i,\alpha}\right)^{2},$
@@
where $B$ denotes the number of batches, $N$ the number of atoms in the batch, $E_{b}$ the ground-truth energy, $\hat{E}_{b}$ the predicted energy, and $F_{i,\alpha}$ the force on atom $i$ in the direction $\alpha\in\{\hat{x},\hat{y},\hat{z}\}$. The weights $\lambda_{E}$ and $\lambda_{F}$ are set to 1 and 100 (for $\Delta^{2}$, 50), respectively.
@@
The models were trained using the AMSGrad variant65 of the Adam optimizer,66 with the default parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. We set the initial learning rate to 0.01 and used a batch size of 5. To adjust the learning rate, we applied an on-plateau scheduler based on the validation loss, with a patience of 500 epochs and a decay factor of 0.9. Additionally, we used an exponential moving average with a weight of 0.99 for evaluation on the validation set and for the final model.
@@
Figure 1: Overall schematic for this work. (a) Architecture of the NequIP-BAM model. (b) Global pooling block: ReLU activation enables learning of the phaseless coupling term $\Delta^{2}$(top), while the identity function shows discontinuities (bottom). (c) Nonadiabatic molecular dynamics schematic
@@
The NequIP-BAM model architecture consists of four major blocks: embedding block, interaction block, self-interaction block, and global pooling block. The embedding block maps atomic chemical properties and positional information into vector space, enabling efficient processing by the model. The interaction block models inter-atomic interactions while maintaining rotational equivariance. The self-interaction block updates features by applying identical weights for each atom and uses different weights according to rotational orders to maintain equivariance. Finally, the global pooling block aggregates features from each atom to generate final predictions. In particular, the global pooling block effectively learns the phaseless coupling term $\Delta^{2}$ by applying ReLU activation function, ensuring positive values. The trained model enables nonadiabatic molecular dynamics that can clearly distinguish between $trans$-PSB3 and $cis$-PSB3 isomers, accurately predicting their positions and dynamical behavior on the potential energy surfaces

### v2

# Machine Learning Nonadiabatic Dynamics: Eliminating Phase Freedom of Nonadiabatic Couplings with the State-Interaction State-Averaged Spin-Restricted Ensemble-Referenced Kohn-Sham Approach

@@ Abstract
Excited-state molecular dynamics (ESMD) simulations near conical intersections (CIs) pose significant challenges when using machine learning potentials (MLPs). Although MLPs have gained recognition for their integration into mixed quantum-classical (MQC) methods, such as trajectory surface hopping (TSH), and their capacity to model correlated electron-nuclear dynamics efficiently, difficulties persist in managing nonadiabatic dynamics. Specifically, singularities at CIs and double-valued coupling elements result in discontinuities that disrupt the smoothness of predictive functions. Partial solutions have been provided by learning diabatic Hamiltonians with phaseless loss functions to these challenges. However, a definitive method for addressing the discontinuities caused by CIs and double-valued coupling elements has yet to be developed. Here, we introduce the phaseless coupling term, $\Delta^{2}$, derived from the square of the off-diagonal elements of the diabatic Hamiltonian in the state-interaction state-averaged spin-restricted ensemble-referenced Kohn-Sham (SI-SA-REKS, briefly SSR)(2,2) formalism. This approach improves the stability and accuracy of the MLP model by addressing the issues arising from CI singularities and double-valued coupling functions. We apply this method to the penta-2,4-dieniminium cation (PSB3), demonstrating its effectiveness in improving MLP training for ML-based nonadiabatic dynamics. Our results show that the $\Delta^{2}$ based ML-ESMD method can reproduce ab initio ESMD simulations, underscoring its potential and efficiency for broader applications, particularly in large-scale and long-timescale ESMD simulations.
@@ 1 Introduction
From a practical perspective, mixed quantum-classical (MQC) approaches17, 18, 19, 20, 21 are commonly used where MQC methods treat electronic propagation quantum mechanically while the nuclei move classically along multiple trajectories. Since conventional photochemical reactions occur over timescales ranging from hundreds of femtoseconds to nanoseconds, a large number of electronic structure calculations are required during an ESMD simulation. For instance, predicting PESs for 100,000 nuclear configurations is necessary to simulate dynamics for 1 ps with a time step size of 1 fs and 100 nuclear trajectories. Moreover, ESMD is highly sensitive to the accuracy of nonadiabatic coupling vectors (NACVs), particularly in cases involving conical intersections (CIs). In such scenarios, quantum chemical methods that adequately account for electronic correlation must be chosen. Although there are highly accurate quantum mechanical methods, e.g. equation-of-motion coupled-cluster singles and doubles method and its derivatives22, 23, 24, 25, second-order perturbation theory with a complete active space self-consistent field26, and density matrix renormalization group27, MQC simulations require hundreds or even thousands of trajectories, making it computationally prohibitive to use such high-level ab initio theories, even for small systems. This makes the trade-off between accuracy and computational cost a central issue in MQC dynamics, presenting an ongoing challenge in balancing accuracy with feasibility.
@@
Over the past decade, machine learning potentials (MLPs) have enabled us to overcome the accuracy-cost trade-off in theoretical chemistry. MLPs can predict molecular properties with an accuracy comparable to traditional ab initio theories but at a significantly lower computational cost. 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50 These advances in ML-based property prediction can be integrated into the MQC dynamics by replacing quantum mechanical calculations with MLPs. However, applying MLPs to ESMD introduces unique challenges that are not typically encountered in ground state MD simulations. A key difficulty is the prediction of NACVs near CIs of PESs. The NACV formula incorporates the energy difference between two electronic states in the denominator. Even minor energy errors can cause significant NACV inaccuracies near CIs. Additionally, NACVs are not differentiable and diverge at CIs. Various studies have proposed solutions to the singularities at CIs. The first approach is to avoid the direct prediction of NACV by using a method that does not require NACV. The simplicity of the Landau–Zener formalism51 has attracted researchers and has been employed in various studies52, 53. The Zhu-Nakamura method54, based on the Landau–Zener formalism, has also been utilized in ML-based noadiabatic dyanmics55, 56. Another approach focused on learning diabatic Hamiltonian instead of adiabatic Hamiltonian57, 58, 59. This provided smooth and differentiable predictions for energy and NACVs near CIs.
@@
While the issue of CI singularity can be addressed through diabatic transformation, another challenge arises from the double-valued nature of the coupling elements. This double-valued nature, known as the geometric phase effect60, imparts a similar double-valued character to NACVs. As a result, NACVs derived from electronic structure calculations have arbitrary signs, leading to discontinuities that hinder efficient model training. To circumvent this problem, a phaseless loss function was introduced61 and applied in several authors of this work62. However, this approach does not fully resolve the problem, as the training set contains two values (one positive and one negative) for the same molecular geometry. Recent advancements in MLPs have introduced various strategies to address the sign problem in ESMD simulations. For adiabatic states, MLPs have been trained to directly predict adiabatic energies while addressing nonadiabatic coupling vectors through techniques such as phaseless loss functions61 or auxiliary single-valued functions63, which circumvent the sign ambiguity inherent in nonadiabatic couplings. In diabatic states, where the sign problem manifests in the off-diagonal elements of the Hamiltonian, similar approaches have been employed, including phaseless loss functions62 to mitigate this issue. Our approach differs by directly training MLPs on phaseless quantities within the diabatic framework, ensuring that the training process avoids the complexities introduced by phase dependence. By leveraging the inherent strengths of MLPs, which utilize both function values and their derivatives for higher accuracy, we can efficiently learn diabatic Hamiltonian matrix elements while avoiding the challenges associated with formulating non-adiabatic coupling vectors in adiabatic states.
@@
Here, we introduce a state-interaction state-averaged spin-restricted ensemble-referenced Kohn-Sham (SI-SA-REKS, briefly SSR)(2,2)64, 65, 66 formalism in combination with training of phaseless quantities to address both the CI singularities and the double value problem of coupling elements. Our approach is based on deriving the SSR(2,2) equations using the phaseless coupling, $\Delta^{2}$, which represents the square of the off-diagonal elements in the diabatic Hamiltonian, rather than the phase-dependent $\Delta$. We trained equivariant message-passing MLPs to model diabatic PESs and the phaseless $\Delta^{2}$ for the penta-2,4-dieniminium cation (PSB3). This method resolves the phase ambiguity of coupling elements and enhances the accuracy of MLP prediction of NACV in SSR(2,2) ESMD simulations.
@@ 2 Method
2.1 SSR(2,2) methodology
We employ the SSR(2,2) methodology to calculate excited electronic states. In the SSR(2,2) method, an ensemble of ground and excited microstates provides $(2\times 2)$ diabatic Hamiltonian, incorporating both perfectly spin-paired singlet (PPS) and open-shell singlet (OSS) configurations. These configurations are determined by two active Kohn-Sham (KS) orbitals, $\phi_{a}$ and $\phi_{b}$, along with their fractional occupation numbers (FONs), $n_{a}$ and $n_{b}$. From these orbitals, a total of six microstates can be configured, enabling the construction of both PPS and OSS configurations as
@@ 2.1 SSR(2,2) methodology
$f(n_{a},n_{b})=(n_{a}n_{b})^{(1-(1/2)((n_{a}n_{b}+\delta)/(1+\delta)))}$ with $\delta=$0.4. State-averaged energy functional $E^{SA}$ is constructed from $E^{PPS}$ and $E^{OSS}$, by $E^{SA}=w_{PPS}E^{PPS}+w_{OSS}E^{OSS}$. Here, we choose equiensemble condition, $w_{PPS}=w_{OSS}=1/2$. From the optimization of orbitals, we obtain the pseudo-Fock equation,
@@
$\hat{F}_{p}=\sum\limits_{L}C^{SA}_{L}{\frac{n^{L}_{p\alpha}\hat{F}^{L}_{\alpha}+n^{L}_{p\beta}\hat{F}^{L}_{\beta}}{n_{p}}}.$ (4)
Here, $n_{p}$ is an averaged occupation $n_{p}=\sum\limits_{L}C_{L}^{SA}(n^{L}_{p\alpha}+n^{L}_{p\beta})$, $n_{p\alpha/\beta}^{L}$ is an occupation number of the $p$th orbital in the $L$the microstate configuration, $C_{L}^{SA}$ is an averaged weight $C_{L}^{SA}=w_{PPS}C_{L}^{PPS}+w_{OSS}C_{L}^{OSS}$, and $\epsilon^{SA}_{pq}$ is a Lagrange matrix element. Finally, we can construct $(2\times 2)$ secular equation to obtain ground and first excited adiabatic state energies, $E_{-}^{SSR}$ and $E_{+}^{SSR}$, as
@@
where we focus on real-valued $\Delta^{SA}$ since conventional quantum chemistry program provides real-valued coupling terms.
The state interaction between PPS and OSS configuration, $\Delta^{SA}$, is calculated from a Lagrange matrix element between two active orbitals $\phi_{a}$ and $\phi_{b}$, and FONs as
$\Delta^{SA}=(\sqrt{n_{a}}-\sqrt{n_{b}})\epsilon^{SA}_{ab}$ (6)
where FONs $n_{a}$ and $n_{b}$ can be obtained from the optimization of KS orbitals while minimizing $E^{SA}$. In the SSR(2,2) method, the NACV between two adiabatic states can be directly calculated from the diabatic elements $E^{PPS}$,$E^{OSS}$, and $\Delta^{SA}$ by
$d_{01}=\frac{1}{E_{+}^{SSR}-E_{-}^{SSR}}((a_{00}a_{01}-a_{10}a_{11})\vec{g}+(a_{00}a_{11}+a_{01}a_{10})\vec{h})$ (7)
where $\vec{g}=\frac{1}{2}\nabla(E^{PPS}-E^{OSS})$ and $\vec{h}=\nabla\Delta^{SA}$ while the coefficients $a_{ij}$ can be expressed with $E^{PPS}$, $E^{OSS}$, and $\Delta^{SA}$ by solving the secular equation.
2.2 Adiabatic energies and NACVs from phaseless quantities
Our key observation is that the NACV of Equation (7) can be rewritten based on the phaseless quantity $\Delta^{2}$ (where $\Delta=\Delta^{SA}$, for simplicity) as follows:
$d_{01}=\frac{nm}{E_{+}^{SSR}-E_{-}^{SSR}}((-\Delta^{2})\nabla(E^{PPS}-E^{OSS})+\frac{1}{2}(E^{PPS}-E^{OSS})\nabla\Delta^{2})$ (8)
@@ 2.2 Adiabatic energies and NACVs from phaseless quantities
$n=\pm 1/\sqrt{(E_{-}-E^{OSS})^{2}+\Delta^{2}}$ and $m=\pm 1/\sqrt{(E_{+}-E^{OSS})^{2}+\Delta^{2}}$, respectively (See the derivation in Supporting Information S1). The presence of the $\pm$ sign in $n$ and $m$ indicates the phase of the NACV, which arises from the phase dependency of the eigenvector in Equation (5). The significance of Eq. 8 lies in the complete substitution of $\Delta$ with $\Delta^{2}$. By this substitution, instead of learning the NACV affected by phase arbitrariness, we focus on learning the components of the NACV that are free from phase arbitrariness of data set which is critical for building MLPs for ESMD. As in the conventional ESMD simulations, the phase of NACVs can be chosen by comparing the sign of NACVs in the previous time step during ESMD simulations. We note that the adiabatic state energy and gradient can be rewritten by $\Delta^{2}$ instead of $\Delta$ as
@@
$\nabla E_{-}^{SSR}$ $=$ $n^{2}((E_{-}^{SSR}-E^{OSS})^{2}\nabla E^{PPS}+\Delta^{2}\nabla E^{OSS}+(E_{-}^{SSR}-E^{OSS})\nabla\Delta^{2})$
$\nabla E_{+}^{SSR}$ $=$ $m^{2}((E_{+}^{SSR}-E^{OSS})^{2}\nabla E^{PPS}+\Delta^{2}\nabla E^{OSS}+(E_{+}^{SSR}-E^{OSS})\nabla\Delta^{2}).$ (10)
In this study, we employed the SSR(2,2) approach with two electronic states only as an application of our machine learning scheme since currently only SSR(2,2) can offer analytic gradients for NAMD simulations. A two-state Hamiltonian effectively describes various photochemical reactions involving nonadiabatic coupling between the S0 and S1 states67, 68, 69, 70, but it can be extended to a three-state Hamiltonian through the training of phaseless functions (see Supporting Information S2 for details).
2.3 Machine learning potential for phaseless quantities
@@
We employed the equivariant message-passing MLP, NequIP33, implemented in the Bayesian Atoms Modeling (BAM) package, to train three key properties: $E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$. Although a single MLP could theoretically model all properties simultaneously, task interference, where gradients from different tasks conflict, usually hinders learning71. And for this reason, we trained three independent MLPs for these diabatic properties, assuming that each property is best represented as a bijective mapping between molecular geometries and output. In training process, the analytical gradients of $\Delta^{2}$, denoted as $\nabla\Delta^{2}$, were constructed as $\nabla\Delta^{2}=2\Delta\nabla\Delta$ from $\Delta$ and $\nabla\Delta$ in the data set, which is also phaseless. To ensure that $\Delta^{2}$ remained positive, we applied inductive bias layers, such as ReLU activation function, enforcing a non-negative value constraint. During the MLP prediction for $\Delta^{2}$, a small shift correction of $5\times 10^{-5}$ eV2 was applied which arose primarily because the ReLU activation function constrains $\Delta^{2}$ predictions to non-negative values (including zero), making it numerically unfeasible to predict exactly zero.
A database was generated consisting of 48,750 data points (40,000, 4,375, and 4,375 geometries for training, validation, and test, respectively), randomly selected from 50 trajectories obtained using exact factorization-based surface hopping dynamics (SHXF)21. These trajectories were computed using the SSR(2,2) formalism combined with the $\omega$PBEh72/6-31G*73 level of theory, with a time step of 0.24 fs over a total simulation duration of 300 fs. Models were trained on a single NVIDIA H100 GPU, with the number of training epochs of 10,000. Averaged training time for three MLPs is 278 GPU-hours. Each model was built with 5 interaction layers and 64 uncoupled feature channels, with $l_{\text{max}}=2$ (Supporting Information S3 and S4). In e3nn notation,74 this is represented as “64x0o + 64x0e + 64x1o + 64x1e + 64x2o + 64x2e”. For all models, radial features were generated using 8 Bessel basis functions and a polynomial envelope with $p=2$ to handle the interatomic interaction within the cutoff.75 These radial features are fed to a multi-layer perceptron $\mathcal{M}$ with layer sizes [64, 64, 64, 1152], using SiLU activation functions in the hidden layers. After the interaction layers, node energies or coupling elements were predicted using a single-layer $\mathcal{M}$ with 16 hidden dimensions. A 6 Å  cutoff was applied to all molecules. The loss function used in training is defined as:
@@
$\mathcal{L}=\frac{\lambda_{E}}{B}\sum_{b=1}^{B}\left(\hat{E}_{b}-E_{b}\right)^{2}+\frac{\lambda_{F}}{3BN}\sum_{i=1}^{B\cdot N}\sum_{\alpha=1}^{3}\left(-\frac{\partial\hat{E}}{\partial r_{i,\alpha}}-F_{i,\alpha}\right)^{2},$ (11)
@@
where $B$ denotes the number of batches, $N$ the number of atoms in the batch, $E_{b}$ the target value ($E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$), $\hat{E}_{b}$ the predicted value, and $F_{i,\alpha}$ the target minus gradient on atom $i$ in the direction $\alpha\in\{\hat{x},\hat{y},\hat{z}\}$. The weights $\lambda_{E}$ and $\lambda_{F}$ were optimized to 1 and 100 (for $\Delta^{2}$, 50), respectively.
@@
The models were trained using the AMSGrad variant76 of the Adam optimizer,77 with the default parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$ (Supporting Information S5). We set the initial learning rate to 0.01 and used a batch size of 5. To adjust the learning rate, we applied an on-plateau scheduler based on the validation loss, with a patience of 500 epochs and a decay factor of 0.9. Additionally, we used an exponential moving average with a weight of 0.99 for evaluation on the validation set and for the final model.
@@
Figure 1: Overall schematic for this work. (a) Architecture of the NequIP-BAM model. (b) Global pooling block: ReLU activation enables learning of the phaseless coupling term $\Delta^{2}$(bottom), while the identity function shows discontinuities (top). (c) Nonadiabatic molecular dynamics schematic
@@
The NequIP-BAM model architecture consists of four major blocks: embedding block, interaction block, self-interaction block, and global pooling block (Figure 1(a)). The embedding block maps atomic chemical properties and positional information into vector space, enabling efficient processing by the model. The interaction block models inter-atomic interactions while maintaining rotational equivariance. The self-interaction block updates features by applying identical weights for each atom and uses different weights according to rotational orders to maintain equivariance. Finally, the global pooling block aggregates features from each atom to generate final predictions. In particular, the global pooling block effectively learns the phaseless coupling term $\Delta^{2}$ by applying ReLU activation function, ensuring positive values (Figure 1(b)). The trained model enables nonadiabatic molecular dynamics that can clearly distinguish between $trans$-PSB3 and $cis$-PSB3 isomers, accurately predicting their positions and dynamical behavior on the potential energy surfaces (Figure 1(c)).